OpenStudy (anonymous):

Simplify: –2m3 • (3m)2 A. –6m2 B. –6m3 C. –18m2 D. –18m5

OpenStudy (mathstudent55):

Why did you close the previous post where you asked this question where I started helping you? Are you interested in learning this or do you just want answers?

OpenStudy (anonymous):

@mathstudent55 Calm down.

OpenStudy (anonymous):

i am intrested in learning i closed it by a accident sorry :(

OpenStudy (mathstudent55):

Ok, no problem. If you want we can continue here, or if you prefer, we can go back to the closed one and continue there.

OpenStudy (anonymous):

we can continue here.... i guess

OpenStudy (mathstudent55):

As long as we're here, let's just continue here. Look at the second part of the expression: \( (3m)^2 \) When you have a product raised to a power, raise every factor to the power. The rule is \( (ab)^m = a^mb^m \) For example, \( (4a)^3 = 4^3 a^3 \) Can you do that to the the part \( (3m)^2 \)?

OpenStudy (anonymous):

no

OpenStudy (mathstudent55):

Each factor inside the parentheses has to be raised to the power, so \( (3m)^2 = 3^2m^2\) Ok?

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

Now let's look at the entire question: \(-2m^3 \cdot (3m)^2 \) We already know what happens to the (3m)^2 part: \(=-2m^3 \cdot 3^2m^2 \) 3^2 is simply 9: \(= -2m^3 \cdot 9m^2\) Now we multiply the numbers together. -2 * 9 = -18 When you multiply the same variable raised to two powers, you write the variable and add the powers. \(m^3 \cdot m^2 = m^{3 + 2} = m^5 \) \(= -18 m^5\) That's the final answer.

OpenStudy (anonymous):

i get it now

OpenStudy (mathstudent55):

Great.